3] A force of 6 N and another of 8 N can be applied together to produce the effect of a single force of:

(a) 1 N (b) 11N

(c) 15 N (d) 20N

Grade:11

4 Answers

vikas askiitian expert
13 years ago

maximum value of net force = Fmax = A+B = 8+6=14N

minimum value of net force = Fmin = A-B = 8-6 = 2

         14 = >  F > = 2

force can take value bw 2 & 14 , so option b) is correct....

Aniket Patra
13 years ago

The maximum value would be 8+6=14N and the minimum would be 8-6=2N.

DEVKANT
4 years ago
given,
1st force=6N and,
2nd force=8N then
maxmium value of net force =(8+6)N=14N and,
minimum value of net force =(8-6)N=2N
so,the effect of single force produced is between 2 to 14N. so answer is 11N(b)
Yash Chourasiya
3 years ago
Hello Student

Here,
Maximum value of Net force = Fmax ​= A+B
= 8 + 6 =14N
Minimum value of Net force = Fmin ​= A−B
= 8 − 2 = 2N
∴ Force can take value as14>F>2
Thus,11N might be the force.

I hope this answer will help you.